Trap6a6yNino44:

Kaylib’s eye-level height is 48 ft above sea level, and Addison’s eye-level height is 85 and one-third ft above sea level. How much farther can Addison see to the horizon? Use the formula d = StartRoot StartFraction 3 h Over 2 EndFraction EndRoot, with d being the distance they can see in miles and h being their eye-level height in feet. StartRoot 2 EndRoot mi 2 StartRoot 2 EndRoot mi 14 StartRoot 2 EndRoot mi 28 StartRoot 2 EndRoot mi

supie:

Can you take a screenshot please?

Trap6a6yNino44:

Trap6a6yNino44:

I think its A

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trap6a6yNino44 I think its A \(\color{#0cbb34}{\text{End of Quote}}\) Not quite,

supie:

brb

Trap6a6yNino44:

Me either but I try

Trap6a6yNino44:

ok

supie:

Ok so remember that we're using the formula \(d=\sqrt{\frac{3h}{2}}\)

Trap6a6yNino44:

yea

supie:

\(d=\sqrt{\frac{3h}{2}}\) \(d=\sqrt{\frac{48}{2}}\) \(d=\sqrt{72}\) \(d=\sqrt{\frac{3h}{2}}\) \(d=\sqrt{{}{2}}\) Yes?

Trap6a6yNino44:

Thanks

supie:

That's for Kaylib's eye level height

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trap6a6yNino44 Thanks \(\color{#0cbb34}{\text{End of Quote}}\) thats not the answer

Trap6a6yNino44:

its not

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trap6a6yNino44 its not \(\color{#0cbb34}{\text{End of Quote}}\) No, it's not.

supie:

We still have Addison's

Trap6a6yNino44:

ok

supie:

Then we get both of them and subtract

Trap6a6yNino44:

okay did det

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trap6a6yNino44 okay did det \(\color{#0cbb34}{\text{End of Quote}}\) Ok, what did you get?

Trap6a6yNino44:

C

Trap6a6yNino44:

Its right